# State-observation augmented diffusion model for nonlinear assimilation with unknown dynamics

### v2

@@ Abstract
keywords:
data assimilation , score-based generative model , Bayesian inference
@@ 1 Introduction
$\begin{cases}\bm{x}_{k}=\mathcal{M}_{k}(\bm{x}_{k-1},\bm{\eta}_{k}),\quad\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X},\\ \bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+\bm{\varepsilon}_{k},\quad\bm{\varepsilon}_{k}\sim\mathscr{D}_{k}^{Y},\end{cases}$ (1)
where $\bm{x}_{k}$ and $\bm{y}_{k}$ denote the state and observation variables at time step $k$, respectively. The observation noise $\bm{\varepsilon}_{k}\sim\mathscr{D}_{k}^{Y}$ captures uncertainties in the observational data, while the term $\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X}$ accounts for possible model errors in the physical model $\mathcal{M}_{k}$. The goal is to estimate the posterior distribution $p(\bm{x}_{S:T}\mid\bm{y}_{S:T})$ for the state variables over the interval from $S$ to $T$, i.e., we use the abbreviation $\bm{x}_{S:T}$ for $\{\bm{x}_{k}\}_{k=S}^{T}$. By Bayes’ rule, the posterior of the states given the observations can be expressed as
@@
$\bm{x}_{S:T}^{*}=\operatorname*{argmin}_{\bm{x}_{S:T}}\left\{\frac{1}{2}\left\|\bm{x}_{S}-\bm{x}_{S}^{\mathrm{b}}\right\|_{\bm{B}}^{2}+\frac{1}{2}\sum_{k=S}^{T-1}\left\|\bm{x}_{k+1}-\mathcal{M}_{k}^{X}(\bm{x}_{k})\right\|_{\bm{Q}_{k}}^{2}+\frac{1}{2}\sum_{k=S}^{T}\left\|\bm{y}_{k}-\mathcal{H}_{k}(\bm{x}_{k})\right\|_{\bm{R}_{k}}^{2}\right\}$
@@
$\bm{x}_{S}\sim\mathcal{N}(\bm{x}_{S}^{\mathrm{b}},\bm{B}),\quad\bm{\eta}_{k}\sim\mathcal{N}(\bm{0},\bm{Q}_{k}),\quad\bm{\varepsilon}_{k}\sim\mathcal{N}(\bm{0},\bm{R}_{k}),$ (5)
@@ 2.1 Data assimilation frameworks
Data assimilation has been studied for decades, and numerous methods have been developed to address the assimilation problem. Broadly, they can be categorized by how they handle the posterior distribution. Variational methods (3D/4D-Var) typically assume Gaussian errors and solve for the MAP estimate, providing a “best-fit” solution. Ensemble-based methods such as the ensemble Kalman filter (EnKF) and other Kalman filter (KF) variants also adopt Gaussian assumptions but use an ensemble of samples to represent the posterior’s mean and covariance. Particle filters, specially designed for nonlinear problem, approximate the posterior by a weighted sum of Dirac deltas, allowing non-Gaussian distributions. Despite recent progress in the particle flow filter (PFF) [29, 30] and its demonstrated potential in high-dimensional settings [31], The need for a large number of particles remains a computational bottleneck. All of these methods rely on physical knowledge and can be prohibitively expensive in high-dimensional problems that require repeated ensemble updates or large-scale optimizations. Despite these challenges, they remain indispensable for reliably integrating observations with complex dynamical models in many scientific and engineering applications.
@@
Nowadays, deep learning (DL)-enhanced data assimilation has gained increasing attention [32]. A pioneering example is the framework of [33], which replaces physical models with neural networks, demonstrating the feasibility of data-driven assimilation. Leveraging the flexibility and fast inference of neural networks, some studies have proposed using them as fast emulators [34] or correctors [35] for physical models in hybrid assimilation frameworks built on classical algorithms. Another research direction focuses on addressing the high-dimensionality challenge through latent assimilation (LA) [36], which seeks a lower-dimensional representation of the state variables to be assimilated using either linear Reduced-Order Models (ROMs) [37, 38] or neural networks [39, 40]. Although methods like LAINR [41] extend the flexibility of LA, they still rely on well-established classical assimilation algorithms as backbones to operate in the latent space.
@@
To better capture non-Gaussian posteriors, recent work has explored the Score-based Filter (SF) [42] and its variants, such as the Ensemble Score-based Filter (EnSF) [43] and Latent-EnSF [44], which rely on diffusion probabilistic models (DPMs) for posterior estimation. These methods have shown promise for nonlinear assimilation tasks but typically assume availability of a physical model. More recently, Score-based Data Assimilation (SDA) [45, 46] has been introduced to integrate physical models into the learning of background priors within a variational assimilation framework, producing promising results for subsampled observations even without any explicit knowledge of physical models. However, as demonstrated in our experiments, SDA still struggles with nonlinear observations. Table 1 summarizes key features of some typical data assimilation approaches. Bolded entries highlight capabilities that surpass those of other methods in the same category.
Method physical model posterior modeling computational cost
3D/4D-Var required MAP high
@@
Table 1: Comparison of some typical data assimilation approaches.
@@ 2.2 Deep generative models
Inspired by [54], the most recent diffusion probabilistic models (DPMs) [55, 56, 57] have achieved impressive performances across various tasks such as image/video generation, restoration, and inpainting [58, 59, 60, 61]. DPMs consider the data generation process as a reversed procedure of a forward diffusion process sending clear states to Gaussian noise. To sample $\bm{x}\sim\mathscr{D}$ from a target distribution, a denoising model is used to recover the unperturbed states from randomly generated Gaussian noise by solving a backward stochastic differential equation [62]. Thus, DPMs need to learn the score function defined as the gradient of the marginal log-likelihood $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ for each time step $t$ during training. In the context of conditional data generation, the gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}\mid\bm{y})$ can be decomposed into the sum of the prior score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ and the gradient of the time-dependent likelihood $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}\mid\bm{x}_{t})$ by Bayes’ rule, and a DPM should focus on establishing surrogate models for both terms. Some works [63, 64] directly replace the gradient term $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}\mid\bm{x}_{t})$ by a neural network. Alternatively, other works like DPS [65] and DMPS [66] have proposed deterministic models in a zero-shot setting for lower computational cost and better performance.
@@ 2.3 Main contributions
In this study, we propose a novel data-driven deep learning (DL) approach specifically designed for posterior estimation to handle the assimilation problem. To address the challenges of nonlinearity, we introduce our State-Observation Augmented Diffusion (SOAD) model, which relaxes the linear constraints typically imposed on physical and observational models in most existing assimilation frameworks. Similar to previous DL-based assimilation methods, our SOAD model is fully data-driven and does not require any prior knowledge of the physical models. Additionally, the state-observation augmented structure not only enables us to assimilate observations from multiple sources simultaneously but also drops the requirement of physical knowledge of observations. The flexibility and adaptability make our approach suitable for the real-world applications. Our main contributions are as follows:
@@ 3.1 Problem settings and assumptions
$\begin{cases}\bm{x}_{k}=\mathcal{M}_{k}^{X}(\bm{x}_{k-1},\bm{\eta}_{k}),\quad\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X},\\ \bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+\bm{\varepsilon}_{k},\quad\bm{\varepsilon}_{k}\sim\mathscr{D}_{k}^{Y},\end{cases}$ (6)
where the additional superscript “$X$” indicates that the model $\mathcal{M}_{k}^{X}$ acts in the physical space for $\bm{x}_{k}$. We follow the setup described in [45, 62], assuming access to a sufficiently large dataset of physical states corresponding to an unknown physical model $\mathcal{M}_{k}^{X}$. All observational operators their noise distributions $\mathscr{D}_{k}^{Y}$ are assumed known as well. Given a sequence of observations $\bm{y}_{S:T}$, our objective is to effectively estimate the posterior distribution of the state variables $\bm{x}_{S:T}$, denoted by $p\left(\bm{x}_{S:T}\mid\bm{y}_{S:T}\right)$.
@@
    Decomposable observational operators: Although the observational operator $\mathcal{H}_{k}$ may vary with time, there exists a time-independent operator $\mathcal{H}$ such that $\mathcal{H}_{k}=\bm{S}_{k}\circ\mathcal{H}$ for some time-varying linear operator $\bm{S}_{k}$.
We make the first assumption to ensure data-driven methods remain valid after training on sufficient historical data. The second assumption is crucial for our theoretical deduction in later sections. These assumptions are pragmatic in many practical contexts. For instance, numerical weather prediction often assumes daily or seasonal periodicity for the physical model $\mathcal{M}_{k}^{X}$. The observational operator $\mathcal{H}_{k}$ may vary at different time steps due to limitations of observational facilities, yet the set of all possible observational operators typically does not change. Hence, we may set
@@
as a concatenation of all possible observational outputs, where each $\mathcal{H}^{(i)}$ corresponds to one type of observation measured on full domain for $i=1,2,\cdots,K$. This suggests that the variation in observational operators for different time steps can be adequately captured by a subsampling matrix $\bm{S}_{k}$ acting on $\mathcal{H}(\bm{x}_{k})$. While traditional assimilation often presume Gaussian noise distributions $\mathscr{D}^{X}$ and $\mathscr{D}^{Y}$, we will demonstrate later that these assumptions can be relaxed as well.
@@ 3.2.1 Score-based generative models
where $f(t)$ and $g(t)$ represent the drift term and diffusion coefficient, respectively, and $\bm{w}_{t}$ denotes the standard Wiener process. The transition kernel from $\bm{x}_{s}$ to $\bm{x}_{t}$ is then given by
@@
is also a diffusion process [68], where the gradient term $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$, known as the score function, plays a crucial role in guiding the reverse diffusion. Abusing of notation, we still let $\bm{w}_{t}$ denote the reverse-time Wiener process.
Consequently, as long as a good estimation of the score function is obtained for all $t$, we may first sample an initial state $\bm{x}_{1}\sim\mathcal{N}(0,\bm{I})$ and then simulate the reverse-time process starting from $\bm{x}_{1}$ to generate a sample $\bm{x}_{0}$ that approximately follow the distribution $p_{\mathrm{data}}(\bm{x})$. Classical numerical solvers for general SDEs, such as the Euler-Maruyama and stochastic Runge-Kutta methods, as well as ancestral sampling [58, 62], are applicable. Recently, to improve sampling efficiency, the exponential integrator (EI) discretization scheme [69]
for better performances.
@@
$=\frac{1}{2}\operatorname{\mathbb{E}}_{\bm{x}_{t}\sim p_{t}}\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right\|^{2}+\operatorname{Tr}\nabla_{\bm{x}_{t}}\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$
$=\frac{1}{2}\operatorname{\mathbb{E}}_{\bm{x}_{0}\sim p_{0}}\operatorname{\mathbb{E}}_{\bm{x}_{t}\sim p_{t\mid 0}(\cdot\mid\bm{x}_{0})}\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right\|^{2}+\operatorname{Tr}\nabla_{\bm{x}_{t}}\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$

### v3

@@ 1 Introduction
$\begin{cases}\bm{x}_{k}=\mathcal{M}_{k}(\bm{x}_{k-1},\bm{\eta}_{k}),\quad\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X},\\ \bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+{\bm{\epsilon}_{k}},\quad{\bm{\epsilon}_{k}}\sim\mathscr{D}_{k}^{Y},\end{cases}$ (1)
where $\bm{x}_{k}$ and $\bm{y}_{k}$ denote the state and observation variables at time step $k$, respectively. The observation noise ${\bm{\epsilon}_{k}}\sim\mathscr{D}_{k}^{Y}$ captures uncertainties in the observational data, while the term $\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X}$ accounts for possible model errors in the physical model $\mathcal{M}_{k}$. The goal is to estimate the posterior distribution $p(\bm{x}_{S:T}\mid\bm{y}_{S:T})$ for the state variables over the interval from $S$ to $T$, i.e., we use the abbreviation $\bm{x}_{S:T}$ for $\{\bm{x}_{k}\}_{k=S}^{T}$. By Bayes’ rule, the posterior of the states given the observations can be expressed as
@@
$\bm{x}_{S:T}^{*}=\operatorname*{argmin}\limits_{\bm{x}_{S:T}}\left\{\frac{1}{2}\left\|\bm{x}_{S}-\bm{x}_{S}^{\mathrm{b}}\right\|_{\bm{B}}^{2}+\frac{1}{2}\sum\limits_{k=S}^{T-1}\left\|\bm{x}_{k+1}-\mathcal{M}_{k}^{X}(\bm{x}_{k})\right\|_{\bm{Q}_{k}}^{2}+\frac{1}{2}\sum\limits_{k=S}^{T}\left\|\bm{y}_{k}-\mathcal{H}_{k}(\bm{x}_{k})\right\|_{\bm{R}_{k}}^{2}\right\}$
@@
$\bm{x}_{S}\sim\mathcal{N}(\bm{x}_{S}^{\mathrm{b}},\bm{B}),\quad\bm{\eta}_{k}\sim\mathcal{N}(\bm{0},\bm{Q}_{k}),\quad{\bm{\epsilon}_{k}}\sim\mathcal{N}(\bm{0},\bm{R}_{k}),$ (5)
@@ 2.1 Data assimilation frameworks
Data assimilation has been studied for decades. In general, data assimilation can be categorized into filtering and smoothing. Filtering aims to estimate the current state given observations up to the current time, while smoothing involves estimating states using both past and future observations. The choice between the two types depends on not only the specific application and objective of the assimilation task but also the availability of data and computational resources.
@@
Ensemble Kalman filter (EnKF) and its variants are representative classical methods for the filtering problem, which adopt Gaussian assumptions and use an ensemble of samples to represent the posterior’s mean and covariance to seek a balance between accurate probabilistic modeling and computational efficiency. In contrast, the 3D-Var method solves for the MAP estimate for the posterior, providing a “best-fit” solution. Particle filters, specially designed for the nonlinear problem, approximate the posterior by a weighted sum of Dirac deltas, allowing non-Gaussian distributions. Despite recent progress in the particle flow filter (PFF) [29, 30] and its demonstrated potential in high-dimensional settings [31], the need for a large number of particles remains a computational bottleneck.
To address the smoothing problem, one of the most intuitive approaches is the Kalman smoother (KS), which includes an additional backward pass to refine estimates using future information compared to the original KF. As ensemble-based extensions of KS, Ensemble Kalman Smoother (EnKS) [32] and iterative EnKS (IEnKS) [33] utilize an ensemble forward propagation to obtain better convergence for nonlinear problems. 4D-Var formulates a variational problem over a time window to retrieve the most likely trajectory, making it inherently a smoothing approach as well.
@@
All of these classical methods rely on physical knowledge and can be prohibitively expensive in high-dimensional problems that require repeated ensemble updates or large-scale optimizations. Despite these challenges, they remain indispensable for reliably integrating observations with complex dynamical models in many scientific and engineering applications. Nowadays, deep learning (DL)-enhanced data assimilation has gained increasing attention [34]. A pioneering example is the framework of [35], which replaces physical models with neural networks, demonstrating the feasibility of data-driven assimilation. Leveraging the flexibility and fast inference of neural networks, some studies have proposed using them as fast emulators [36] or correctors [37] for physical models in hybrid assimilation frameworks built on classical algorithms. Another research direction focuses on addressing the high-dimensionality challenge through latent assimilation (LA) [38], which seeks a lower-dimensional representation of the state variables to be assimilated using either linear Reduced-Order Models (ROMs) [39, 40] or neural networks [41, 42]. Although methods like LAINR [43] extend the flexibility of LA, they still rely on well-established classical assimilation algorithms as backbones to operate in the latent space.
@@
To better capture non-Gaussian posteriors, recent work has explored the Score-based Filter (SF) [44] and its variants, such as the Ensemble Score-based Filter (EnSF) [45] and Latent-EnSF [46], which rely on diffusion probabilistic models (DPMs) for posterior estimation. These methods have shown promise for nonlinear assimilation tasks but typically assume availability of a physical model. More recently, Score-based Data Assimilation (SDA) [47, 48] has been introduced to integrate physical models into the learning of background priors within a variational assimilation framework, producing promising results for subsampled observations even without any explicit knowledge of physical models. However, as demonstrated in our experiments, SDA still struggles with nonlinear observations.
In the modern numerical weather prediction pipeline, increasing attention has been given to smoothing-based methods, particularly 4D-Var and its variants [49, 50, 51, 52], due to their ability to incorporate time-evolving observations and improve the dynamical consistency of the atmospheric state [53, 54]. These advances highlight a growing need for learning-based smoothing algorithms that can effectively assimilate observations over a time window and handle the non-Gaussianity arising from nonlinear dynamics and observation operators. Motivated by this, our proposed method adopts a smoothing perspective and seeks to estimate latent trajectories conditioned on both past and future observations, with a focus on addressing challenges related to nonlinearity and posterior complexity. Table 1 summarizes key features of some typical data assimilation approaches. Bolded entries highlight capabilities that surpass those of other methods in the same category.
Method physical model posterior modeling comput. cost
@@
4D-Var required MAP high
KS, EnKS[32], IEnKS[33] required mean & covariance high
@@
Table 1: Comparison of some typical filtering and smoothing approaches.
@@ 2.2 Deep generative models
Inspired by [62], the most recent diffusion probabilistic models (DPMs) [63, 64, 65] have achieved impressive performances across numerous data generation and restoration tasks, particularly in computer vision and natural language processing domains. Specifically, DPMs have shown state-of-the-art performances in image generation and restoration tasks such as denoising, inpainting, and super-resolution [66, 67, 68], as well as in text-to-image and video generation [69]. These tasks often involve linear measurement operators (e.g., masking, blurring, or downsampling), and thus the associated inverse problems are typically linear, enabling efficient deployment of diffusion models through techniques like iterative reconstruction guided by the learned score function or posterior sampling [70].
More recently, diffusion models have been adapted to tackle complex inverse problems in scientific computing. Unlike many standard image restoration scenarios, scientific computing tasks such as phase retrieval [71, 72], seismic inversion [73, 74, 75], and topology optimization [76, 77, 78] generally involve nonlinear, high-dimensional, and often ill-posed inverse problems. In these contexts, DPMs are leveraged to model the prior distribution of states efficiently and sample from complex posterior distributions conditioned on nonlinear observations.
@@ 2.3 Main contributions
In this study, we propose a novel data-driven deep learning (DL) approach specifically designed for posterior estimation to handle the assimilation problem. To address the challenges of nonlinearity, we introduce our State-Observation Augmented Diffusion (SOAD) model, which relaxes the linear constraints typically imposed on physical and observational models in most existing assimilation frameworks. Similar to previous DL-based assimilation methods, our SOAD model is fully data-driven and does not require any prior knowledge of the physical models. Importantly, while our network training still relies on observational data that implicitly reflects the observation process, our model does not require an explicit physical formulation of the observational operator. That is, as long as paired state-observation samples are available, SOAD can be trained and subsequently perform assimilation with observations from multiple heterogeneous sources without requiring any knowledge of how these observations are physically related to the state variables. The flexibility and adaptability make our approach suitable for the real-world applications. Our main contributions are as follows:
@@ 3.1 Problem settings and assumptions
$\begin{cases}\bm{x}_{k}=\mathcal{M}_{k}^{X}(\bm{x}_{k-1},\bm{\eta}_{k}),\quad\bm{\eta}_{k}\sim\mathscr{D}_{k}^{X},\\ \bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+{\bm{\epsilon}_{k}},\quad{\bm{\epsilon}_{k}}\sim\mathscr{D}_{k}^{Y},\end{cases}$ (6)
where the additional superscript “$X$” indicates that the model $\mathcal{M}_{k}^{X}$ acts in the physical space for $\bm{x}_{k}$. We follow the setup described in [47, 79], assuming access to a sufficiently large dataset of physical states corresponding to an unknown physical model $\mathcal{M}_{k}^{X}$. The noise distributions $\mathscr{D}_{k}^{Y}$ associated with all observational operators are assumed to be known as well. Given a sequence of observations $\bm{y}_{S:T}$, our objective is to effectively estimate the posterior distribution of the state variables $\bm{x}_{S:T}$, denoted by $p\left(\bm{x}_{S:T}\mid\bm{y}_{S:T}\right)$.
@@
    Decomposable observational operators: Although the observational operator $\mathcal{H}_{k}$ may vary with time, there exists a time-independent operator $\mathcal{H}$ such that $\mathcal{H}_{k}=\bm{S}_{k}\circ\mathcal{H}$ for some time-varying linear operator $\bm{S}_{k}$ with orthogonal rows.
These assumptions are pragmatic in many practical contexts. For instance, numerical weather prediction frequently exhibits daily or seasonal periodicity in the underlying physical model $\mathcal{M}_{k}^{X}$. The first assumption ensures the continued applicability and effectiveness of data-driven methods trained on historical data, a standard practice in deep learning-based assimilation methodologies [47, 36, 41, 39, 43]. The observational operator $\mathcal{H}_{k}$ typically varies over time due to observational constraints or infrastructure limitations [80], but the range of possible observational configurations usually remains fixed. Therefore, we may set
@@
as a concatenation of all possible observational outputs, where each $\mathcal{H}^{(i)}$ corresponds to one type of observation measured on full domain for $i=1,2,\cdots,K$. This suggests that the variation in observational operators for different time steps can be adequately captured by a “subsampling matrix” $\bm{S}_{k}$ acting on $\mathcal{H}(\bm{x}_{k})$. Here, a subsampling matrix is a binary matrix that selects a subset of elements (or locations) from the full state vector. While traditional assimilation often presume Gaussian noise distributions $\mathscr{D}^{X}$ and $\mathscr{D}^{Y}$, we will demonstrate later that these assumptions can be relaxed as well.
@@ 3.2.1 Score-based generative models
where $f(t)$ and $g(t)$ represent the drift term and diffusion coefficient, respectively, and $\bm{w}_{t}$ denotes the standard Wiener process. The transition kernel from $\bm{x}_{0}$ to $\bm{x}_{t}$ is then given by
@@
is also a diffusion process [82], where the gradient term $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$, known as the score function, plays a crucial role in guiding the reverse diffusion. With a slight abuse of notation, we still let $\bm{w}_{t}$ denote the reverse-time Wiener process. Consequently, as long as a good estimation of the score function is obtained for all $t$, we may first sample an initial state $\bm{x}_{1}\sim\mathcal{N}(0,\bm{I})$ and then simulate the reverse-time process starting from $\bm{x}_{1}$ to generate a sample $\bm{x}_{0}$ that approximately follow the distribution $p_{\mathrm{data}}(\bm{x})$.
@@
$=\frac{1}{2}\operatorname{\mathbb{E}}_{\bm{x}_{t}\sim p_{t}}{\left[\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right\|^{2}+\operatorname{Tr}\nabla_{\bm{x}_{t}}\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right]}$
$=\frac{1}{2}\operatorname{\mathbb{E}}_{\bm{x}_{0}\sim p_{0}}\operatorname{\mathbb{E}}_{\bm{x}_{t}\sim p_{t\mid 0}(\cdot\mid\bm{x}_{0})}{\left[\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right\|^{2}+\operatorname{Tr}\nabla_{\bm{x}_{t}}\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)\right]}$
@@
Once a good estimator for the score funtion is obtained, it suffices to evolve the backward diffusion process by discretizing the time steps. Classical numerical solvers for general SDEs, such as the Euler-Maruyama and stochastic Runge-Kutta methods, as well as ancestral sampling [66, 79], are applicable. Recently, to improve sampling efficiency, the exponential integrator (EI) discretization scheme [86]
for better performances, where we let $\delta$ denote the time step.